OpenStudy (zenmo):

Find the orthogonal trajectories of the family of curves. (Use C for any needed constant.) Question on answer, last step.

OpenStudy (zenmo):

\[y^2=7kx^3\]

OpenStudy (zenmo):

How should I begin this problem?

OpenStudy (wavybaby):

Where's C?

OpenStudy (zenmo):

C will probably end up in the end-product, but I think K is the constant?

OpenStudy (wavybaby):

So you trying to solve the problem by finding k?

OpenStudy (wavybaby):

I would tell you what i got but, there's no answer choices and i'm not a 100% sure. Don't wanna get you wrong. ;)

OpenStudy (zenmo):

\[\frac{ y^2 }{ 7x^3 }=k\]Using quotient rule: \[\frac{ 14x^3yy'-21x^2y^2 }{ 49x^6 }=0\]\[\frac{ 2xyy'-3y^2 }{ 7x^4 }=0\]\[xyy'-3y^2=0, 2xy'-3y=0, 2xy'=3y, y' = \frac{ 3y }{ 2x }=> \frac{ -dx }{ dy }= \frac{ 3y }{ 2x }\]\[-2xdx = 3ydy\]\[\int\limits_{}^{}-2xdx = \int\limits_{}^{}3ydy\] I am at the last step, what shall I do now?

OpenStudy (zenmo):

Okay, if I do it the normal way, the answer is\[-x^2+C = \frac{ 3 }{ 2 }y^2, C = x^2+\frac{ 3 }{ 2 }y^2\] Another slightly different approach, the book's way, \[-x^2+\frac{ C }{ 2 }=\frac{ 3y^ 2}{ 2 }, -2x^2 + C = 3y^2, 2x^2+3y^2=C\] this is also an answer. I am wondering why it used C/2 instead of just "C"?

OpenStudy (photon336):

\[\frac{ dy }{ dx } => \frac{ 3y^{2} }{ 2xy }\] how did you get -dx/dy?